### v1

# LLaVA-MoD: Making LLaVA Tiny via MoE-Knowledge Distillation

Abstract
We introduce LLaVA-MoD, a novel framework designed to enable the efficient training of small-scale Multimodal Language Models (s-MLLM) by distilling knowledge from large-scale MLLM (l-MLLM). Our approach tackles two fundamental challenges in MLLM distillation. First, we optimize the network structure of s-MLLM by integrating a sparse Mixture of Experts (MoE) architecture into the language model, striking a balance between computational efficiency and model expressiveness. Second, we propose a progressive knowledge transfer strategy to ensure comprehensive knowledge migration. This strategy begins with mimic distillation, where we minimize the Kullback-Leibler (KL) divergence between output distributions to enable the student model to emulate the teacher network’s understanding. Following this, we introduce preference distillation via Direct Preference Optimization (DPO), where the key lies in treating l-MLLM as the reference model. During this phase, the s-MLLM’s ability to discriminate between superior and inferior examples is significantly enhanced beyond l-MLLM, leading to a better student that surpasses its teacher, particularly in hallucination benchmarks. Extensive experiments demonstrate that LLaVA-MoD outperforms existing models across various multimodal benchmarks while maintaining a minimal number of activated parameters and low computational costs. Remarkably, LLaVA-MoD, with only 2B activated parameters, surpasses Qwen-VL-Chat-7B by an average of 8.8% across benchmarks, using merely $0.3\%$ of the training data and 23% trainable parameters. These results underscore LLaVA-MoD’s ability to effectively distill comprehensive knowledge from its teacher model, paving the way for the development of more efficient MLLMs. The code will be available on https://github.com/shufangxun/LLaVA-MoD
Figure 1: Comparisons of training cost and performance. LLaVA-MoD achieves competitive performance with advanced MLLMs using significantly lower training costs while outperforming current small-scale MLLMs by a large margin.
1 Introduction
Leveraging the advanced instruction-following and reasoning capabilities of Large Language Models (LLMs) (Achiam et al., 2023; Bai et al., 2023a; Jiang et al., 2024; Team et al., 2023; Touvron et al., 2023a; b; Dubey et al., 2024), Multimodal Large Language Models (MLLMs) (Bai et al., 2023b; Liu et al., 2024; Lin et al., 2024b; Li et al., 2023b; Chen et al., 2023b; Shu et al., 2023; Lu et al., 2024), typically integrate a visual encoder (Caron et al., 2021; Radford et al., 2021; Liu et al., 2022) alongside an LLM to achieve promising results across various vision and cross-modal tasks, such as image captioning and visual question-answering. MLLMs are characterized by their large model size and extensive training data, which contribute to their superior performance but also necessitate significant computation resources. For instance, the largest version of LLaVA-NeXT (Li et al., 2024a) employs Qwen-1.5-110B (Yang et al., 2024) as the language model and requires 18 hours of training with 128 H800 GPUs. Furthermore, MLLMs face deployment challenges due to the high memory and computational demands. For instance, models with substantial parameters often require advanced hardware and exhibit slow inference speeds, which significantly hinder their real-world applications such as on mobile devices. Thus, developing MLLMs that balance performance and efficiency has become a crucial research focus.
Previous works on small-scale Multimodal Large Language Models (s-MLLM) (Zhou et al., 2024a; Yuan et al., 2023; Shao et al., 2024; He et al., 2024; Chu et al., 2023; 2024; Yao et al., 2024) have mainly concentrated on devising data collection (Schuhmann et al., 2021; 2022; Zhu et al., 2024; Awadalla et al., 2024; Gadre et al., 2024; Chen et al., 2023a) and filtering protocols (Xu et al., 2023; Fan et al., 2024; Zhang et al., 2023; Li et al., 2024b; Nguyen et al., 2024) to ensure high-quality training data. While such high-quality data aid s-MLLM in narrowing the performance gap with their larger counterparts, their effectiveness is still limited by model capacity and the availability of high-quality data. With the advent of open-source MLLMs, an intuitive strategy inspired by conventional deep learning tasks involves utilizing large-scale MLLMs (l-MLLM) as the teacher to improve the training of s-MLLM via Knowledge Distillation (KD) (Hinton et al., 2015). KD facilitates stable and efficient training of smaller models by aligning the output distributions of the student model with those of the teacher model. However, the application of KD as a model reduction technique has yet to be extensively explored in MLLMs. To develop an effective framework for MLLM distillation, we consider two primary challenges. The first challenge lies in designing a lightweight architecture for student MLLMs that retains strong learning and expressive capabilities, enabling it to absorb the complex knowledge embedded within teacher MLLMs effectively. The second challenge is to efficiently and comprehensively transfer this knowledge from the teacher MLLMs to the student MLLMs.
To address these two challenges, we present LLaVA-MoD [^1] MoD denotes Mixture-of-Expert Knowledge Distillation for MLLM, an effective framework for training s-MLLM by mimicking the behavior of the l-MLLM through Mixture-of-Expert (MoE) Knowledge Distillation. For the first challenge—designing a lightweight s-MLLM structure—a straightforward approach might involve reducing the scale of the base language model in l-MLLM to create a smaller network. However, this direct reduction in network parameters can significantly compromise the model’s expressive capability, making it less effective in modeling complex multimodal understanding tasks. Drawing inspiration from the recent success of sparse MoE (Lin et al., 2024a; Dai et al., 2024; Jiang et al., 2024; Shen et al., 2023) in sequential modeling, we incorporate a sparse MoE structure into the dense s-MLLM. This approach seeks to balance scale reduction while preserving the model’s ability to capture and represent intricate multimodal information during distillation. Specifically, we enhance the s-MLLM by equipping it with several feedforward networks (FFNs) and linear gates within the LLM. Each FFN acts as an expert, capturing fine-grained knowledge from the l-MLLM, while the gates select the top-$k$ experts to facilitate optimal knowledge transfer pathways.
To address the second challenge, we propose a progressive distillation strategy for effective knowledge transfer. The process begins by aligning the vision encoder with the LLM using a learnable adapter, thereby initializing a dense student MLLM. Following this, we employ two consecutive distillation stages, where the student model evolves from mimicking and approximating the teacher MLLM to ultimately surpassing it: Mimic Distillation. This stage is divided into two steps, i.e., dense-to-dense and dense-to-sparse distillation. In the dense-to-dense distillation stage, we employ standard knowledge distillation loss to align the output logits distribution between the initialized dense s-MLLM and the teacher model, using general captioning and conversation datasets. Next, we transition to the dense-to-sparse distillation where we transform the dense student MLLM into a sparse one by integrating MoE, and then distill knowledge from the teacher MLLM into this sparse student model using a broad range of tasks and datasets. Preference Distillation. In this stage, the teacher model provides knowledge regarding what constitutes “good” and “bad” samples, establishing a foundational reference for the student model. The student MLLM leverages this knowledge to adjust its probability distribution, ensuring that good samples have a higher probability than those from the teacher model, while bad samples are assigned a lower probability. This process enhances the student model’s ability to mitigate hallucinations by improving its judgment capabilities beyond those of the teacher model.
As illustrated in Figure 1, LLaVA-MoD exhibits impressive performance across various multimodal benchmarks while maintaining minimal activated parameters and low computational resources. For example, LLaVA-MoD-2B exceeds Qwen-VL-Chat-7B by an average of $8.8\%$ on these benchmarks, utilizing just $0.3\%$ of the training data and $23\%$ trainable parameters. Furthermore, it matches the performance of RLHF-based methods with 7B and 13B parameters on several hallucination benchmarks. Specifically, LLaVA-MoD-2B surpasses RLHF-V (Yu et al., 2024a) by $8.2\%$ in response-level hallucination rate and by $21.3\%$ in mention-level hallucination rate on the Object HalBench. The impressive results demonstrate the effectiveness of our proposed MoD framework in transferring knowledge from l-MLLM to s-MLLM.
2 Related Work
Multimodal Large Language Models. The emergence of large language models (LLMs) has greatly advanced the field of natural language processing. Connecting visual information to LLM to enhance their understanding of multimodal input is crucial for promoting the unified comprehension of vision and language. CLIP (Radford et al., 2021) was the first to align visual and textual information into a unified embedding space through contrastive learning objectives. BLIP-2 (Li et al., 2023b) utilizes various pre-training tasks to train separate visual and text streams, and adds additional intermediary structures to adapt visual features to LLM. Flamingo (Alayrac et al., 2022) incorporates additional cross-attention modules into LLM to handle arbitrarily interleaved sequences of images and texts. Recently, models like LLaVA (Liu et al., 2024) and MiniGPT-4 (Zhu et al., 2023) use linear projection layers to map image features to the textual space and enhance the models’ instruction-following ability through visual instruction tuning. Moreover, some works focus on a stronger vision encoder. Intern-VL (Chen et al., 2023b) leverages a larger-scale vision encoder to bridge the gap with a large-scale language model, and Mini-Gemini (Li et al., 2024c) introduces additional vision encoders for high-resolution refinement to enhance visual tokens. Qwen-VL (Bai et al., 2023b), VisionLLM (Wang et al., 2023c), and others aim to imbue models with stronger fine-grained visual understanding capabilities, such as grounding and region understanding. Unlike these approaches, our method does not emphasize enlarging the model to boost multimodal capabilities. Instead, we adopt a distillation approach to compress the capabilities of larger models into a smaller, sparse MoE architecture, enhancing computational and storage efficiency.
Knowledge Distillation. Large language models possess powerful capabilities, but their enormous size and high inference costs limit their application in low-resource environments. Knowledge distillation (Hinton et al., 2015) uses a large model as teacher to transfer its advanced knowledge to a smaller student model, which plays a critical role in compressing model size and enables smaller models to self-improve. Some works adopt additional designs to make the distillation process more suitable for LLMs. MiniLLM (Gu et al., 2023) minimizes the reverse KLD (Kullback–Leibler divergence) to prevent students from overestimating low-probability regions in the teacher distribution, while GKD (Agarwal et al., 2023) introduces generalized knowledge distillation and promotes the integration of distillation with RLHF (Reinforcement Learning from Human Feedback). Additionally, Huang et al. (Huang et al., 2022) introduce context distillation, and some works (Mukherjee et al., 2023; Li et al., 2022; Ho et al., 2023) adopt CoT (Chain-of-Thought) distillation to enhance specific skills of small models. Our approach innovatively distills the knowledge of MLLM into a smaller sparse MoE (Mixture-of-Experts) architecture, significantly enhancing the multimodal processing capabilities of small models at a minimal cost.
Mixture-of-Experts The mixture of experts (MoE) architecture is initially introduced by Jacobs et al. (1991) to enhance model performance across various samples through the utilization of independent experts. Recently, with the rise of transformer architectures, numerous studies have extended transformer layers to incorporate MoE. In transformer-based MoE architecture, feed-forward neural network (FFN) layers are often replaced with sparsely activated experts, employing a trainable Top-k gating strategy (Lepikhin et al., 2020; Fedus et al., 2022), thereby effectively increasing model capacity while maintaining a low computational overhead. Additionally, to further mitigate training costs, the sparse upcycling method (Komatsuzaki et al., 2022) is adopted, which initializes expert parameters using those from a well-trained dense model. Currently, MoE architectures have found widespread application not only in pre-trained language models (Jiang et al., 2024; Dai et al., 2024) but also in vision models and vision-language models (Lin et al., 2024a; Shen et al., 2023). Our approach integrates MoE with knowledge distillation techniques to provide stronger signals for sparse training, which remarkably decrease the training cost associated with sparse models.
3 Method
Figure 2: Progressive Distillation of LLaVA-MoD. (a). Mimic Distillation: Aligning the student’s response probabilities $p_{S}$ with those of the teacher $p_{T}$, via the Kullback–Leibler (KL) Loss. (b). Preference Distillation: Increasing the student’s positive response probabilities $p^{+}_{S}$ to surpass those of the teacher $p^{+}_{T}$, while decreasing the student’s negative response probabilities $p^{-}_{S}$ to fall below those of the teacher $p^{-}_{T}$, via the Preference-Optimization (PO) Loss.
We introduce LLaVA-MoD, a novel framework for building efficient MLLM using mixture-of-experts (MoE) knowledge distillation. Our framework consists of two main components: (a). Architecture Design of s-MLLM: As shown in Figure 3, we design the s-MLLM with a sparse MoE framework, enhancing the ability to acquire specialized expert knowledge. (b). Distillation Mechanism: We design a progressive distillation mechanism as shown in Figure 2 to transfer knowledge from the l-MLLM to the sparse s-MLLM. This process involves two stages: mimic distillation followed by preference distillation. We will introduce LLaVA-MoD in detail from these two components.
3.1 Architecture Design of Sparse s-MLLM
In this section, we describe the architecture design of our sparse s-MLLM, which serves as the student model in the distillation process.
s-MLLM Definition.
As illustrated in Figure 3, the basic architecture of s-MLLM consists of three primary components: a vision encoder, a large language model (LLM), and a vision-language (VL) adaptor. Given a multimodal instruction conversation $(x,y)$, we define our s-MLLM to process response $y$ as follows:
$y=\text{LLM}_{\phi}(\text{Proj}_{\omega}(\text{ViT}_{\chi}(x_{v})),x_{i}),$ (1)
where $x_{v}$ is the input image, and $x_{i}$ is the text instruction. The input image is resized to 336 $\times$ 336 and patched into 576 image tokens, each of size 14 $\times$ 14. $\text{ViT}_{\chi}$ is the CLIP vision encoder with parameters $\chi$ which first extracts image features from $x_{v}$. $\text{Proj}_{\omega}$ is the vision-language adaptor with parameters $\omega$, serving as the vision tokenizer to align the image features with the word embedding space. $\text{LLM}_{\phi}$ is the large language model with parameters $\phi$, which produces the response $y$ based on the multimodal tokens of $x=[x_{v},x_{i}]$.
Figure 3: Dense to Sparse Architecture of LLaVA-MoD. The architecture of LLaVA-MoD consists of three components: VL Adaptor, Vision Encoder, and LLM. The VL Adaptor and Vision Encoder remain unchanged, while the LLM is upcycled from dense to sparse.
Mixture-of-Experts.
The principle of building our s-MLLM is downsizing the LLM while leaving the vision encoder and vision-language adaptor unmodified. To achieve this downsizing goal, we sparsify the dense s-MLLM by incorporating an MoE architecture. Specifically, Figure 3 illustrates the process, where we apply the sparse upcycling technique (Komatsuzaki et al., 2022) to replicate $N$ feedforward networks (FFNs) as the expert modules. Additionally, we introduce a linear layer as the router, which dynamically activates the appropriate experts by predicting the probability of expert assignment. Given each token $x$ in the sequence, we first compute the routing value of $N$ experts:
$r=\text{Softmax}(x\cdot W_{r}),$ (2)
where $W_{r}$ denotes the weight matrix of the router and each element $r_{i}$ in $r$ represents the probability of activating the $i$-th expert. After that, we apply the Top-$k$ strategy to determine the activated experts with the highest $k$ routing values:
$\tilde{r}=\text{Top-}k({r})=\begin{cases}r_{i},&\text{if }i\in k\\ 0,&\text{otherwise}\\ \end{cases}$ (3)
where the routing values for the inactivated experts are set to zero, effectively excluding them from contributing to the final output. The output $y$ is calculated by aggregating the contributions of the activated experts, weighted by their corresponding routing values:
$y=\sum_{i=1}^{N}\tilde{r}\cdot E_{i}(x).$ (4)
3.2 Progressive Distillation
Our progressive distillation consists of two distinct stages, i.e., mimic distillation (Figure 2 (a)) and preference distillation (Figure 2 (b)). In the mimic distillation stage, the student MLLM $\pi_{\text{S}}$ imitates the general and specific knowledge from the teacher MLLM $\pi_{\text{T}}$. In the preference distillation stage, the student MLLM gains the preference knowledge of the teacher MLLM to further refine its output and reduce hallucinations. We use the pre-trained l-MLLM as the teacher MLLM, which is frozen during the whole distillation process. Both teacher and student MLLMs are in the same LLM family, ensuring consistent vocabulary space, which is essential for the student MLLM to imitate the teacher MLLM accurately. For simplicity, we denote the student as $\pi_{\text{S}}$ and the teacher as $\pi_{\text{T}}$.
Initialization.
Before distillation, we first align the vision encoder with the LLM via a learnable adapter, aiming to obtain a well-initialized dense student MLLM. The $\text{LLM}_{\phi}$ and $\text{ViT}_{\chi}$ are maintained frozen since their pre-trained parameters have already captured rich visual and language knowledge. Only $\text{Proj}_{\omega}$ is optimized to bridge the gap between the vision and language domain. For the initialization, we utilize common image-caption pairs from a widely used and curated dataset, which covers a diverse range of subjects and visual entities. The training objective is to minimize the cross-entropy of the generated tokens. The objective function is defined as:
$\mathcal{L}_{\text{Init}}(\pi_{\text{S}})=-\mathbb{E}_{(y_{k}\mid y_{<k},x)\sim\pi_{\text{S}}}\left[\log\pi_{\text{S}}(y_{k}\mid y_{<k},x)\right],$ (5)
where $\pi_{\text{S}}(y_{k}\mid y_{<k},x)$ represents the probability of the predicted token $y_{k}$ conditioned on $x$ and the previous tokens $y_{<k}=(y_{1},y_{2},\dots,y_{k-1})$.
Mimic Distillation.
We decompose the comprehensive knowledge of the teacher MLLM into general and expertise aspects and undertake dense-to-dense distillation and dense-to-sparse distillation respectively to transfer this knowledge to the student MLLM.
a). Dense-to-Dense Distillation. In this stage, we aim to replicate the general knowledge of the teacher MLLM. Acquiring general knowledge first is crucial, as it provides a broad foundation and common understanding across various domains, enabling the student MLLM to develop a basic framework applicable in multiple scenarios. This foundation supports a more comprehensive and flexible understanding before progressing to specialized tasks. To achieve this, we maintain $\text{ViT}_{\chi}$ frozen, and jointly optimize $\text{LLM}_{\phi}$ and $\text{Proj}_{\omega}$, with trainable parameters $\theta=\{\omega,\phi\}$. We leverage common image-caption pairs and conversation datasets. The training objective is to minimize the Kullback-Leibler divergence (KLD) between the output logits of the dense student MLLM and the teacher MLLM. The objective function is defined as:
$\mathcal{L}_{\text{D2D}}(\pi_{\text{S}};\pi_{\text{T}})=-\mathbb{E}_{(x,y_{k})\sim\pi_{\text{T}}}\left[\log\frac{\pi_{\text{T}}(y_{k}\mid y_{<k},x)}{\pi_{\text{S}}(y_{k}\mid y_{<k},x)}\right],$ (6)
where $V$ denotes the vocabulary, while $\pi_{\text{S}}(y_{k}\mid y_{<k},x)$ and $\pi_{\text{T}}(y_{k}\mid y_{<k},x)$ denote the probability of the predicted tokens for student and teacher models, respectively.
b). Dense-to-Sparse Distillation. In this stage, our focus shifts to transferring the specialized expertise of the teacher MLLM, allowing the student MLLM to obtain advanced capabilities and achieve superior performance in diverse tasks. However, directly learning this knowledge in its dense form could lead to inefficiencies and difficulty. Therefore, we sparsify the dense student MLLM by introducing multiple experts. As detailed in Section 3.1, we replicate $N$ feedforward networks (FFNs) within $\text{LLM}_{\phi}$ and incorporate an MLP layer as a router, which jointly constitutes the experts with parameters $\phi_{e}$. This sparse architecture allows the student MLLM to selectively activate the most relevant experts for different tasks and inputs, offering significant advantages in emulating the teacher’s specialized knowledge. For training, we leverage multi-task data, updating only the experts and the adaptor. We employ a Top-$k$ routing strategy to select the experts. The trainable parameters are $\theta=\{\omega,\phi_{e}\}$. Similar to the previous stage, we also adopt the KLD as the training objective. Additionally, we include the standard next-token prediction objective, i.e., minimizing the student-generated tokens’ cross-entropy loss, to inject supervision from ground truth data which can reduce the existing bias in the teacher model. The final objective function is defined as:
$\mathcal{L}_{\text{D2S}}(\pi_{\text{S}};\pi_{\text{T}})=-\mathbb{E}_{(x,y_{k})\sim\pi_{\text{S}}}\left[\log\pi_{\text{S}}(y_{k}\mid y_{<k},x)\right]-\mathbb{E}_{(x,y_{k})\sim\pi_{\text{T}}}\left[\log\frac{\pi_{\text{T}}(y_{k}\mid y_{<k},x)}{\pi_{\text{S}}(y_{k}\mid y_{<k},x)}\right].$ (7)
Preference Distillation.
In this stage, our goal is to distill the preference knowledge from the teacher MLLM to guide the student MLLM towards generating not only accurate but also reasonable responses, which is crucial in reducing hallucinations. For the training process, we make effective use of preference data, which comprises meticulously paired positive responses $y^{+}$ and negative responses $y^{-}$ for the identical prompt $x$. Our preference distillation strategy is inspired by recent advancements in Direct Preference Optimization (DPO) (Rafailov et al., 2024), which bypasses the need for training a reward model by directly training on an offline preference dataset. Our key insight is to treat the teacher MLLM as the reference model during distillation. The teacher model plays a crucial role by providing insights into what constitutes “good” and “bad”, thereby establishing a fundamental reference for the student model.
Specifically, the training objective is to optimize the student model to assign higher probabilities to positive responses than those of the teacher model while assigning lower probabilities to negative responses than those of the teacher model. This training process involves two key optimization aspects: First, the student model seeks to align with the teacher model in distinguishing positive from negative responses. Second, the student model aims to surpass the teacher model by assigning higher probabilities to positive responses and lower probabilities to negative responses. We only train the mixture of experts and the VL adaptor in the student model and employ a Top-$k$ routing strategy to select the experts. The trainable parameters are $\theta=\{\omega,\phi_{e}\}$. The object function is defined as:
$\mathcal{L}_{\text{PD}}(\pi_{\text{S}};\pi_{\text{T}})=-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\text{S}}(y^{+}\mid x)}{\pi_{\text{T}}(y^{+}\mid x)}-\beta\log\frac{\pi_{\text{S}}(y^{-}\mid x)}{\pi_{\text{T}}(y^{-}\mid x)}\right)\right],$ (8)
where $\pi_{\text{S}}(y^{+}\mid x)$ and $\pi_{\text{S}}(y^{-}\mid x)$ denote the probabilities of positive and negative responses for the student model, respectively, and $\pi_{\text{T}}(y^{+}\mid x)$ and $\pi_{\text{T}}(y^{-}\mid x)$ denote the same for the teacher model.
4 Experiments
4.1 Experimental Settings
Implementation Details.
We employ the well-established ”ViT-MLP-LLM” architecture to demonstrate the effectiveness of LLaVA-MoD. Specifically, a pre-trained CLIP-ViT-L/14 is utilized as the vision encoder and a 2-layer MLP is utilized as the adaptor for both the student and the teacher models. Both teacher and student models leverage the Qwen-1.5 and Qwen-2 families to construct their base model. Specifically, the teacher model employs a 7B parameter configuration, while the student model is constructed with 1.8B and 0.5B parameter sizes. The performance of teacher MLLM on common multimodal benchmarks is presented in Table 1.
Table 1: Performance of teacher MLLM on multimodal benchmarks.
[…]
Qwen-1.5-7B CLIP-L 62.3 40.7 70.9 55.3 72.1 68.5 64.9 62.1
Qwen-2-7B CLIP-L 62.5 37.9 73.2 57.2 78.0 71.6 71.1 64.5
As introduced in Section 3.2, the training process comprises mimic distillation and preference distillation, each with specific settings. At initialization, we first freeze the vision encoder and LLM while optimizing the VL adaptor to align image tokens with the word embedding space. This stage employs a cross-entropy loss with a batch size of 512 and a learning rate of 1e-4. In the mimic distillation stage, the vision encoder remains frozen, while the LLM and VL adaptor are co-optimized to distill general knowledge from the teacher MLLM in a dense-to-dense fashion. Then, the FFN in the LLM is first transformed into a sparse architecture with a mixture of FFN experts co-optimized with the VL adaptor to distill specialized knowledge from the teacher MLLM in a dense-to-sparse fashion. This stage employs a KL-divergence loss, with cross-entropy loss added for dense-to-sparse distillation. The batch size is 256, and the learning rate is adjusted to 2e-5. In the preference distillation stage, the model inherits from the mimic distillation. The vision encoder remains frozen, and a mixture of FFN experts is co-optimized with the VL adaptor to distill preference knowledge from the teacher MLLM. This stage employs a DPO loss to optimize the student’s probability for positive responses to be greater than that of the teacher and the student’s probability for negative responses to be lower than that of the teacher. The batch size is 256, and the learning rate is adjusted to 2e-6. Throughout all stages, we employ the adam optimizer (Diederik, 2014) and train on 16 NVIDIA A100 GPUs for one epoch each, totaling approximately 960 GPU hours. The detailed hyperparameters of each training stage are illustrated in Appendix  11.
Training Datasets.
The training data consists of 5M samples from the open-source datasets, with each training stage utilizing a distinct dataset. During Initialization, 0.6M general captioning samples from LLaVA-1.5-pretrain dataset (Liu et al., 2023b) are used to bridge the gap between visual and language modalities. In mimic distillation, 2.4M general captioning and conversation samples are used to distill general knowledge from the teacher MLLM and 1.4M multi-tasks data, including VQA, documents, science, and OCR, are used to distill specialized knowledge from the teacher MLLM. For preference distillation, 0.8M preference data samples are used to transfer preference knowledge from the teacher. The detailed dataset of each training stage is illustrated in Appendix 10.
Evaluation Benchmarks.
[…]
4.2 Main Results
In this section, we conduct experiments of LLaVA-MoD to highlight its advantages in two aspects: performance and efficiency. For performance, we evaluate comprehension-oriented benchmarks (Table 2) and hallucination-oriented benchmarks (Table 3). For efficiency, we present a comparison in terms of data samples and model size.
[…]
BLIP-2 Vicuna-13B 129M $\geq$7B 41.0 19.6 61.0 42.5 64.7 - - -
VILA-7B LLaMA-7B 50M 62.3 57.8 68.2 64.4 76.7 68.9 61.7 65.7
CogVLM Vicuna-7B 1500M 64.9 - 65.6 78.2 71.8 63.7 53.8 -
InstructBLIP Vicuna-13B 130M 49.5 33.4 63.1 50.7 60.6 - - -
IDEFICS-80B LLaMA-65B 6.8M 38.4 35.5 59.9 25.9 64.2 48.2 25.2 42.5
Qwen-VL-Chat Qwen-7B 1450M 57.5 38.9 68.2 61.5 74.4 60.6 56.7 56.7
SPHINX-Intern2 InternLM2-7B 15M 56.2 49.6 70.4 58.1 63.2 57.9 - -
Deepseek-VL-7B DLLM-7B 2000M 61.3 49.9 74.0 64.7 73.4 74.1 72.8 67.2
LLaVA-1.5-7B Vicuna-1.5-7B 1.2M 62.0 50.0 66.8 58.2 75.5 64.3 58.3 62.1
LLaVA-NeXT Vicuna-1.5-13B 1.3M 65.4 60.5 73.6 67.1 78.7 70.0 64.4 68.5
Imp-3B Phi-2-2.7B 1.6M $\sim$3B 63.5 54.1 72.8 59.8 72.3 72.9 46.7 63.2
Bunny-3B Phi-2-2.7B 2.7M 62.5 43.8 70.9 56.7 74.4 68.6 37.2 59.2
VILA-3B LLaMA-2.7B 51M 61.5 53.5 69.0 60.4 72.1 63.4 52.7 61.8
LLaVA-Phi Phi-2-2.7B 1.3M - 35.9 68.4 48.6 66.7 59.8 - -
TinyLLaVA Phi-2-2.7B 3.2M 61.0 - 70.1 53.5 71.6 68.3 - -
TinyGPT-V Phi-2-2.7B 24M 33.6 33.4 - - - - - -
MobileVLM MLLaMA-2.7B 1.3M 59.0 - 61.0 47.5 64.4 59.6 - -
MobileVLMv2 MLLaMA-2.7B 3.6M 61.1 - 70.0 57.5 72.0 63.2 - -
MoE-LLaVA-3B Phi-2-2.7B 2.2M 61.4 43.9 68.5 51.4 71.1 65.2 41.8 57.6
MiniCPM-V MiniCPM-2.4B 570M 51.5 50.5 74.4 56.6 68.9 64.0 62.7 61.2
MiniCPM-V-2 MiniCPM-2.4B 570M 52.1 60.2 76.3 73.2 70.5 68.5 67.2 66.9
Imp-2B Qwen-1.5-1.8B 1.6M $\sim$2B 61.9 39.6 66.1 54.5 65.2 63.8 61.2 58.9
Bunny-2B Qwen-1.5-1.8B 2.7M 59.6 34.2 64.6 53.2 65.0 59.1 58.5 56.3
Mini-Gemini-2B Gemma-2B 2.7M 60.7 41.5 63.1 56.2 67.0 59.8 51.3 57.1
MoE-LLaVA-2B Qwen-1.5-1.8B 2.2M 61.5 32.6 63.1 48.0 64.6 59.7 57.3 55.3
DeepSeek-VL-1.3B DLLM-1.3B 2000M 59.3 36.8 64.2 58.4 65.3 64.6 61.0 58.5
LLaVA-MoD-2B Qwen-1.5-1.8B 5M 58.7 39.2 68.0 58.5 66.7 66.3 61.9 59.9
LLaVA-MoD-2B Qwen-2-1.5B 5M 58.8 40.4 69.2 59.9 69.2 68.9 64.4 61.6
SPHINX-Tiny TLLaMA-1.1B 15M $\sim$1B 58.0 49.2 21.5 57.8 63.1 56.6 37.8 49.2
LLaVA-MoD-1B Qwen-1.5-0.5B 5M $\sim$1B 56.2 31.6 62.8 53.9 65.3 58.8 50.4 54.1
LLaVA-MoD-1B Qwen-2-0.5B 5M 56.6 35.1 61.1 57.1 67.0 58.7 54.1 55.7
4.2.1 Comprehension-Oriented Benchmarks
As indicated in Table 2, LLaVA-MoD demonstrates remarkable performance on comprehension-oriented benchmarks. It achieves SOTA average results among the models of 2B size and 1B size. Specifically, The 2B-sized LLaVA-MoD surpasses Mini-Gemini-2B (Li et al., 2024c) by 8.1%, while using a lower image resolution (336 vs. 768). The 1B-sized LLaVA-MoD surpasses SPHINX-Tiny (Gao et al., 2024) by 13.2%, using fewer data samples (5M vs. 15M). Furthermore, LLaVA-MoD-2B matches and even surpasses the performance of large-scale MLLMs. Specifically, The 2B-sized LLaVA-MoD surpasses Qwen-VL-Chat-7B (Bai et al., 2023b) by 8.8% and matches the performance of VILA-3B (Lin et al., 2024b) and MiniCPM-V (Yao et al., 2024). These results highlight that our approach effectively and efficiently training small-scale MLLMs by distilling the sparse MoE architecture.
Table 3: Comparison with state-of-the-art MLLMs on the hallucination benchmarks. We compare LLaVA-MoD with SFT-based works and RLHF-based works . Hall: Hallucination Rate Resp: response-level hallucination rate, Ment: mention-level hallucination rate. The best result is in bold, and the second-best result is underlined.
Model LLM #Param Object POPE MMHal
HalBench Bench
Resp $\downarrow$ Ment $\downarrow$ F1 $\uparrow$ Score $\uparrow$ Hall $\downarrow$
Teacher MLLM Qwen-1.5-7B 7B 50.1 24.8 84.9 2.60 20.7
Teacher MLLM Qwen-2-7B 7B 29.7 23.4 85.7 2.88 14.5
Qwen-VL-Chat Qwen-7B 9.6B 40.4 20.7 74.9 2.76 38.5
LLaVA-1.5-7B Vicuna-7B 7B 53.6 25.2 86.1 2.36 51.0
VCD Vicuna-1.5-7B 7B 48.8 24.3 84.5 2.12 54.2
OPERA Vicuna-1.5-7B 7B 45.1 22.3 85.4 2.15 54.2
HA-DPO Vicuna-1.5-7B 7B 39.9 19.9 86.8 1.98 60.4
POVID Vicuna-1.5-7B 7B 48.1 24.4 86.3 2.08 56.2
LLaVA-RLHF Vicuna-1.5-13B 13B 38.1 18.9 82.7 2.02 62.5
LURE Vicuna-1.5-7B 7B 27.7 17.3 - 1.64 60.4
RLHF-V Vicuna-13B 13B 12.2 7.5 86.2 2.45 51.0
RLAIF-V Vicuna-1.5-7B 7B 8.5 4.3 - 3.06 29.2
MiniCPM-V MiniCPM-2.4B 2.8B 21.6 11.5 79.5 3.70 24.9
MiniCPM-V-2 MiniCPM-2.4B 2.8B 14.5 7.8 86.3 4.09 18.2
Mini-Gemini-2B Gemma-2B 2B 29.7 21.1 85.6 2.83 18.8
Bunny-2B Qwen-1.5-1.8 2.2B 50.2 23.4 85.8 2.72 19.3
LLaVA-MoD-2B Qwen-1.5-1.8B 2.2B 11.4 7.2 87.0 2.76 17.2
LLaVA-MoD-2B Qwen-2-1.5B 1.9B 11.2 5.9 87.2 2.91 13.8
4.2.2 Hallucination-Oriented Benchmarks
As indicated in Table 3, LLaVA-MoD shows remarkable performance in mitigating hallucination, even beating its teacher model. It can be attributed into two aspects: Firstly, by assigning a higher probability for the positive response, preference distillation encourages the student model to focus on providing correct and relevant information. Secondly, by assigning a lower probability for the negative response, preference distillation discourages incorrect or unsubstantiated information. Using the teacher model as a reference to adjust response probabilities, this optimization enables the student model to handle hallucination problems more accurately and reliably, thereby outperforming the teacher model.
Moreover, LLaVA-MoD surpasses similar small-sized models (Yao et al., 2024; Li et al., 2024c; He et al., 2024) and even recent RLHF-based models (Sun et al., 2023; Zhou et al., 2024b; Huang et al., 2024; Leng et al., 2024; Zhou et al., 2023; Zhao et al., 2023b). For instance, on the Object HalBench, it outperforms RLHF-V (Yu et al., 2024a) by 8.2% in response-level hallucination rate and by 21.3% in mention-level hallucination rate. This demonstrates that taking the teacher model as a reference for preference distillation is a challenging task that can encourage the student model to effectively learn to mitigate hallucination.
4.2.3 Efficiency Comparison
We compare the efficiency of LLaVA-MoD with recent SOTA MLLMs in terms of training data and trainable parameters. As indicated in Table 2, The 2B-sized LLaVA-MoD requires only 5M samples and activates merely a subset of its total 2.2B parameters during training and inference. It demonstrates that LLaVA-MoD achieves 8.8% higher performance than Qwen-VL-Chat, using only 23% trainable parameters and 0.3% of training data samples, respectively. Moreover, when compared to the small-sized model MiniCPM-V with 2.8B trainable parameters, LLaVA-MoD also achieves superior performance with just 1.6% of the data samples and 2B parameters, underscoring its efficiency. In conclusion, LLaVA-MoD enables more efficient utilization of data, parameters, and computational resources. This not only results in faster training and inference but also provides a practical solution for deploying high-performance models in resource-constrained environments.
4.3 Ablation Study
In this section, we begin by analyzing the impact of preference distillation. Next, we conduct ablation studies on a model trained solely with mimic distillation to investigate the effectiveness of our proposed architecture and training strategy on comprehension-oriented benchmarks. We utilize CLIP-ViT-L/14 as the default vision encoder. The default large language models (LLMs) are Qwen-1.5-1.8B and Qwen-1.5-7B for the student and teacher models.
4.3.1 Impact of Preference Distillation
We extensively evaluated preference distillation, examining its impact on both the comprehension ability and hallucination mitigation capabilities of LLaVA-MoD. As detailed in Table 4 and Table 5, preference distillation does not yield consistent improvements in comprehension across all benchmarks, its impact on hallucination mitigation was strikingly consistent. Across all evaluated benchmarks, preference distillation notably reduced the occurrence of hallucinations, highlighting its efficacy in enhancing the reliability and factual accuracy of LLaVA-MoD. This observation aligns with the field of language modeling (Achiam et al., 2023), have consistently indicated that RLHF techniques tend to prioritize the reduction of hallucinations, often at the expense of potential gains in other areas, such as raw comprehension scores.
Table 4: Performance of preference distillation on comprehension-oriented benchmarks.
[…]
Mimic 59.3 40.0 68.4 58.7 68.4 65.1 61.5 60.2
Mimic + Preference 58.7 39.2 68.0 58.5 66.7 66.3 61.9 59.9
Table 5: Performance of preference distillation on hallucination-oriented benchmarks.
Distillation Stage Object POPE MMHal
HalBench Bench
Resp $\downarrow$ Ment $\downarrow$ F1 $\uparrow$ Score $\uparrow$ Hall $\downarrow$
Mimic 39.1 22.6 86.7 2.75 17.8
Mimic + Preference 11.4 7.2 87.0 2.76 17.2
4.3.2 Impact of Training Strategy
Knowledge Distillation Improves Training Efficiency in MoE.
This study explores the impact of knowledge distillation (KD) and examines how it facilitates MoE model training compared to supervised fine-tuning (SFT). The experiments are conducted using the same sparse configuration of E4T2, where 4 experts are initialized and the top-2 experts are activated. The same datasets and training hyperparameters are employed for both KD and SFT.
Training MoE models is inherently challenging, as it requires effectively leveraging huge data to optimize experts’ routing. Our approach demonstrates that KD, compared to traditional SFT, excels in addressing this challenge, achieving superior performance across multiple benchmarks. As demonstrated in Table 6, KD significantly outperforms SFT on all the benchmarks. Specifically, KD achieves an average improvement of 8.1% over SFT, with particularly pronounced gains in complex multi-task scenarios. For instance, on the MMB and MME benchmarks, KD surpasses SFT by 8.2% and 10.0%, respectively. These results highlight KD’s ability to effectively leverage multiple experts and extract comprehensive knowledge from large-scale MLLMs. By distilling essential knowledge, KD not only maximizes data utilization but also achieves better generalization ability of MoE compared to SFT.
Table 6: Comparison between the KD and SFT. The architecture of the student model is configured as E4T2.
[…]
SFT 58.6 31.2 66.2 55.6 63.2 59.2 56.1 55.7
KD 59.3 40.0 68.4 58.7 68.4 65.1 61.5 60.2
Focusing on Response Improves Distillation Generalization.
This study explores the impact of the distillation target. The default approach distills response, we experiment with distilling both response and instruction. Both settings utilize the same datasets and training hyperparameters. Table 7 shows that distilling only the response achieves superior performance, surpassing jointly distilling response and instruction by 3.1% in terms of average performance. This suggests that focusing on response is more effective for knowledge distillation in auto-regressive modeling. A possible explanation lies in that mimicking instructions could lead to overfitting the specific instruction patterns of training data, potentially reducing the model’s adaptability to unseen instructions. Mimicking the teacher model’s response proves to be effective in improving the student’s auto-regressive modeling generalization capabilities.
Table 7: Comparison between different distillation tokens. The response indicates distilling solely on the generated response tokens, while the response+instruction incorporates additional instruction tokens in the distillation process.
[…]
Response + Instruction 58.5 35.0 66.8 58.3 68.5 61.8 59.0 58.4
Response 59.3 40.0 68.4 58.7 68.4 65.1 61.5 60.2
4.3.3 Impact of Model Architecture
Sparse MoE Enables Effective Knowledge Distillation.
This study explores the effect of sparse architecture in distillation by comparing its performance to dense architecture. We adopt the E4T1 configuration as the sparse architecture to ensure the same activated parameters as the dense architecture. The same datasets and training hyperparameters are used for sparse and dense architectures.
[…]
Table 8: Comparison between sparse and dense architecture within the distillation. The sparse architecture is configured as E4T1.
[…]
Dense 57.6 32.7 67.3 56.8 65.2 61.8 58.2 57.1
Sparse 58.7 36.9 67.9 58.2 66.1 64.5 61.7 59.2
Teacher Capacity Matters in Knowledge Distillation.
This study explores the impact of the teacher in the distillation process. Specifically, we utilize CLIP-ViT-L/14 as the vision encoder. For the LLM, Qwen-1.5-7B is the strong teacher while Qwen-1.5-4B is the weaker teacher. All experiments are conducted under the same datasets and training hyperparameters to ensure a fair comparison.
As shown in Table 9, a suitable teacher enhances performance across various benchmarks. For instance, the Qwen-1.5-1.8B student model achieves an average performance of 62.3% with the weaker Qwen-1.5-4B teacher. This performance jumps to 63.6% when utilizing the stronger Qwen-1.5-7B teacher. Interestingly, the smaller Qwen-1.5-0.5B student model shows marginal improvement with either the Qwen-1.5-4B or Qwen-1.5-7B teacher. This suggests that an excessively large capacity gap between the teacher and student models can hinder effective knowledge transfer. Utilizing a ”middle teacher” with an intermediate capacity could bridge this gap, facilitating smoother knowledge transfer and boosting the student model’s learning efficiency.
Table 9: Comparison between the strong and weak teachers within the distillation. We set the strong teacher as the LLM is Qwen-1.5-7B and the weak teacher as the LLM is Qwen-1.5-4B. The model configuration is E4T2.
[…]
Qwen-1.5-0.5B Qwen-1.5-4B 56.0 25.3 64.7 53.8 63.3 62.2 50.8 53.7
Qwen-1.5-7B 56.1 29.8 63.8 54.5 64.2 58.5 49.7 53.8
Qwen-1.5-1.8B Qwen-1.5-4B 58.7 34.6 67.9 57.7 67.6 64.9 60.7 58.9
Qwen-1.5-7B 59.3 40.0 68.4 58.7 68.4 65.1 61.5 60.2
5 Conclusion
In this paper, we introduce LLaVA-MoD, a novel framework for efficient training of small-scale multimodal language models via knowledge distillation from large-scale ones. It addresses two key challenges in MLLM distillation: enhancing s-MLLM architecture with MoE design for efficiency and expressiveness balance and implementing a progressive knowledge transfer strategy. Extensive experiments show that LLaVA-MoD outperforms existing models with low activation parameters and computational costs. Notably, it outperforms Qwen-VL-Chat-7B by 8.8% with only 2 billion activation parameters, 0.3% training data, and 23% trainable parameters, highlighting its effectiveness in distilling knowledge and driving more efficient multimodal language model development.
6 Limitations
LLaVA-MoD requires that both the teacher and student models belong to the same family of large language models (LLMs) to ensure consistency in the vocabulary space. Future research could explore distillation techniques involving heterogeneous model families. Additionally, the need to load both student and teacher models results in significant memory usage. To facilitate more efficient distillation, a feasible solution would be to pre-extract the outputs of the teacher model, allowing only the student model to be loaded during training.
Appendix A Appendix
A.1 Dataset
We report the detailed training dataset settings of LLaVA-MoD in Table 10.
Table 10: Training dataset. #Sample means the training samples.
Stage Task Dataset
Initialization Captioning LLaVA-1.5-Pretrain (Liu et al., 2023b)
Dense-to-Dense Distillation Captioning ALLaVA-Caption-4V (Chen et al., 2024), ShareGPT4V-PT (Chen et al., 2023a)
Conversation MIMIC-IT (Li et al., 2023a), LVIS (Wang et al., 2023a), LRV (Liu et al., 2023a), SViT (Zhao et al., 2023a)
Dense-to-Sparse Distillation Captioning ShareGPT4V-100K (Chen et al., 2023a), TextCaps (Sidorov et al., 2020)
Conversation LLaVA-1.5-Instruct (Liu et al., 2023b)
General QA GQA (Hudson & Manning, 2019), VQAv2 (Goyal et al., 2017), OKVQA (Marino et al., 2019)
Grounding Visual Genome (Krishna et al., 2017), RefCOCO (Yu et al., 2016; Mao et al., 2016)
Science AI2D (Kembhavi et al., 2016), ScienceQA (Lu et al., 2022a)
Chart & Doc DVQA (Kafle et al., 2018), ChartQA (Masry et al., 2022), DocQA (Clark & Gardner, 2018)
OCR OCRVQA (Mishra et al., 2019), SynthDoG-EN (Kim et al., 2022)
Knowledge A-OKVQA (Schwenk et al., 2022), GeoQA+ (Cao & Xiao, 2022)
Preference Distillation Preference RLAIF-V (Yu et al., 2024b)
A.2 Hyperparameters
We report the detailed training hyperparameters of LLaVA-MoD in Table 11.
Table 11: Training hyperparameters
Configuration Stage-1 Stage-2 Stage-3
LLM
VL Adaptor
ViT
LLM init. Qwen-1.5-1.8B Qwen-1.5-1.8B 2nd-stage
VL Adaptor init. MLP 1st-stage 2nd-stage
ViT init. CLIP-Large@336
Image resolution 336 $\times$ 336
ViT sequence length 576
LLM sequence length 2048
Optimizer AdamW
Optimizer hyperparameter $\beta_{1}=0.9,\beta_{2}=0.98,eps=1e^{-6}$
Learning rate 1e-4 2e-5 2e-5
Learning rate schedule Cosine decay
Weight decay 0.0
Training epoch 1
Warm-up ratio 0.03
Global batch size 256 128 128
Numerical precision Bfloat16
Model parallelism Zero2 Zero2 offload Zero2 offload